# SPEED: Scalable, Precise, and Efficient Concept Erasure for Diffusion Models

### v1

@@ 3.2 Null-Space Constraints
where $\|\boldsymbol{\Delta}\mathbf{P}\|^{2}$ is an additional regularization term to guarantee convergence. This objective facilitates model parameter editing in the null space of the retain set $\mathbf{R}$, showing great promise for prior-preserved concept erasure.
Figure 3: Prior shift comparison with erasing different target concepts. The prior shift is measured by $\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}\|^{2}$ introduced in Sec. 4.1, where $\boldsymbol{c}$ denotes the concept embedding. A larger value indicates a greater impact on the concept after erasure.
@@ 4 Prior Knowledge Refinement
The core of null-space constraints in SPEED is to determine a projection matrix $\mathbf{P}$ that constrains the solution space of $\boldsymbol{\Delta}$ within the null space of the retain set $\mathbf{R}$, ensuring that non-target concepts remain unaffected after erasure. The rank of the matrix $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ reflects the diversity of the retain set, directly shaping the structure of its null space, which is determined by $\mathbf{R}$. As $\mathbf{R}$ increases, $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ gradually reaches full rank, its null space narrows and reduces to the trivial null space $\{\mathbf{0}\}$. However, shrinking $\mathbf{R}$ would eliminate prior coverage by allowing more unintended shifts in non-target concepts not included. To balance this dilemma, we introduce three complementary techniques: Influence-Based Prior Filtering (IPF) to retain the most affected non-target concepts while avoiding excessive rank saturation, Directed Prior Augmentation (DPA) to expand prior coverage with semantically consistent variations, and Invariant Equality Constraints (IEC) to enforce additional constraints to preserve key invariants in generation.
@@ 4.1 Influence-Based Prior Filtering (IPF)
Given a pre-defined retain set, existing editing-based methods [18, 19] generally neglect to account for the varying influence of different target concepts on non-target elements, where closer semantic relationships (e.g., Mickey and Dog for Snoopy) lead to greater prior shifts, as shown in Fig. 3. This differing influence suggests that not all non-target concepts contribute equally to preserving prior knowledge, and weakly influenced concepts provide marginal benefit while introducing additional ranks to constrain the null space. To improve, we propose to filter the original retain set by selectively retaining those non-target concepts most influenced by the erased target concepts.
To evaluate this influence, we first obtain the closed-form solution $\boldsymbol{\Delta}_{\text{erase}}$, which only considers the erasure efficacy during editing. Specifically, we remove the preservation error $e_{0}$ from Eq. 1 and introduce an additional regularization term $\|\boldsymbol{\Delta}\|^{2}$ to guarantee convergence:
$\boldsymbol{\Delta}_{\text{erase}}$ $=\mathop{\text{arg\ min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\boldsymbol{\Delta}\|^{2}}_{\text{regularization}}$
$=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\mathbf{I}+\mathbf{C}_{1}\mathbf{C}_{1}\right)^{-1}.$ (4)
Then we can quantify the influence on each concept embedding $\boldsymbol{c}$ by measuring the prior shift $=\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}\|^{2}$. This metric stems from the intuition that after erasing the target concept, a larger prior shift in the non-target concept indicates a greater influence on the corresponding prior knowledge. Building upon this, we can filter the original retain set $\mathbf{R}$ with the prior shift by the following process:
$\mathbf{R}_{f}:\mathbf{R}\mapsto\{\boldsymbol{c}_{0}\in\mathbf{R}\mid\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}_{0}\|^{2}>\mu\},$ (5)
where the mean value $\mu=\mathbb{E}_{\boldsymbol{c}_{0}\sim\mathbf{R}}\left[\|\boldsymbol{\Delta}_{\text{erase}}\boldsymbol{c}_{0}\|^{2}\right]$ serves as the filtering threshold.
@@ 4.2 Directed Prior Augmentation (DPA)
To prevent IPF from generating an overly small retain set, which would compromise prior coverage, an intuitive strategy is to augment the prior knowledge by perturbing the non-target embedding $\boldsymbol{c}_{0}$ with random noise [37]. However, this strategy would introduce meaningless embeddings that fail to generate semantically coherent images (e.g., noise image), resulting in excessive preservation with increasing ranks. To search for more semantically consistent concepts, we introduce directed noise by projecting the random noise $\boldsymbol{\epsilon}$ onto the direction in which the model parameters $\mathbf{W}$ exhibit minimal variation. This operation ensures that the perturbed embeddings express closer semantics to the original concept after being mapped by $\mathbf{W}$, as shown in Fig. 4. Specifically, we first derive a projection matrix $\mathbf{P}_{\text{min}}$:
@@
where $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$ denotes the eigenvectors w.r.t. the smallest $r$ singular values, which represent the $r$ least-changing directions of $\mathbf{W}$. Then the directed noise $\boldsymbol{\epsilon}\cdot\mathbf{P}_{\text{min}}$ is used to perturb the original embedding via:
$\boldsymbol{c}^{\prime}_{0}=\boldsymbol{c}_{0}+\boldsymbol{\epsilon}\cdot\mathbf{P}_{\text{min}},\quad\boldsymbol{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (7)
@@
$\mathbf{R}^{\text{aug}}:\mathbf{R}\mapsto\bigcup_{\boldsymbol{c}_{0}\in\mathbf{R}}\left\{\boldsymbol{c}_{0,k}^{\prime}\mid k=1,\dots,N_{A}\right\},$ (8)
where $N_{A}$ denotes the augmentation times and $\boldsymbol{c}_{0,k}^{\prime}$ represents the $k$-th augmented embedding given $\boldsymbol{c}_{0}\in\mathbf{R}$ using Eq. 7. In implementation, we first filter the original retain set $\mathbf{R}$ to obtain $\mathbf{R}_{f}$ using IPF. Subsequently, further augmentation and filtering are applied to $\mathbf{R}_{f}$ using DPA and IPF to obtain $\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$, and the two filtered retain sets are then combined together to serve as the final refined retain set $\mathbf{R}_{\text{refine}}=\mathbf{R}_{f}\cup\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$.
Figure 5: Qualitative comparison of the few-concept erasure in erasing single and multiple instances. The erased and preserved generations are highlighted with red and green boxes, respectively. Our method demonstrates superior prior preservation compared to the SOTA method, UCE [18], in single-concept erasure, and this superiority is amplified as the target concepts increase. For example, in the Hello Kitty case, our generation exhibits less alteration after erasing three concepts, whereas UCE’s generations are gradually deteriorating.
@@ 4.3 Invariant Equality Constraints (IEC)
In parallel, we also identify certain invariants during the T2I generation process, i.e., intermediate variables that remain unchanged with sampling prompts. One such invariant is the CLIP-encoded [SOT] token. Since the encoding process is masked by causal attention and all prompts are prefixed with the fixed [SOT] token during tokenization, the corresponding embedding remains unchanged. Another invariant is the null-text embedding, as it corresponds to the unconditional generation under the classifier-free guidance [23], which also remains unchanged despite prompt variations. Given the invariance of these embeddings, we consider additional protection measures for these invariant embeddings to ensure their outputs remain unchanged during concept erasure. Specifically, we introduce explicit equality constraints for invariants based on Eq. 3.2:
$\begin{gathered}\mathop{\text{min}}\limits_{\boldsymbol{\Delta}}\underbrace{\|(\mathbf{W}+\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\boldsymbol{\Delta}\mathbf{P}\|^{2}}_{\text{regularization}},\\ \quad\text{s.t.}\,\underbrace{(\boldsymbol{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}}_{\text{equality constraints}},\end{gathered}$ (9)
where $\mathbf{C}_{2}$ denotes the stacked invariant embedding matrix of [SOT] and null-textSince the null-text embeddings are only composed of [EOT] tokens (excluding [SOT]), we use the k-means algorithm [38] to select k centroids to reduce redundancy.. With the projection matrix $\mathbf{P}$ derived from $\mathbf{R}_{\text{refine}}$, we can compute the closed-form solution of Eq. 9 using Lagrange Multipliers from Appx. B.3:
$\left(\boldsymbol{\Delta}\mathbf{P}\right)_{\text{SPEED}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\mathbf{P}\mathbf{Q}\mathbf{M},$ (10)
@@
$\begin{gathered}\mathbf{M}=\left(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I}\right)^{-1},\\ \mathbf{Q}=\mathbf{I}-\mathbf{M}\mathbf{C}_{2}\left(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2}\right)^{-1}\mathbf{C}_{2}^{\top}\mathbf{P}.\end{gathered}$ (11)
@@ 5 Experiments
In this section, we conduct extensive experiments on three representative erasure tasks, including few-concept erasure, multi-concept erasure, and implicit concept erasure, validating our superior prior preservation. The compared baselines include ConAbl [29], MACE [35], RECE [19], and UCE [18], which have achieved SOTA performance across various concept erasure tasks. In implementation, we conduct all experiments on SDv1.4 [1] and generate each image using DPM-solver sampler [34] over 20 sampling steps with classifier-free guidance [23] of 7.5. More implementation details and compared baselines (e.g., SPM [37]) can be found in Appx. C and Appx. D.2.
@@ 5.1 On Few-Concept Erasure
Evaluation setup. To compare the few-concept erasure performance with baseline methods, we conduct experiments on instance erasure and artistic style erasure following [37], where all methods are evaluated based on 80 instance templates and 30 artistic style templates, generating 10 images per template per concept. We use two metrics for evaluation: CLIP Score (CS) [45] measuring the text-image similarity and Fréchet Inception Distance (FID) [22] assessing the distributional distance before and after erasure. Following [37], we select non-target concepts with similar semantics to the target concept for comparison and report CS for targets and FID for non-targets in the main paper. Full comparisons are presented in Appx. D.1. We further compare the generations on MS-COCO captions [33], where we generate images with the first 1,000 captions, and report the CS and FID as the measure of general knowledge preservation.
@@
Table 1: Quantitative comparison of the few-concept erasure in erasing instances, where the erased target concept(s) are shaded in pink. Arrows on the headers indicate the preferred direction for each metric, and the best results are highlighted in bold. Our method, which does not achieve the lowest CS but has been proven sufficient in Fig. 5, consistently demonstrates significant improvements in prior preservation for non-target concepts.
Analysis and discussion. Tables 1 and 2 compare the results of erasing various instance concepts and artistic styles. Our method consistently achieves the lowest FIDs across all non-target concepts, demonstrating superior prior preservation with minimal alteration to the original content. Moreover, we emphasize that our erasure is sufficiently effective, even without achieving the lowest CS, as shown in Fig. 5. In contrast, lower CS values typically indicate over-erasure, which results in excessive degradation of prior knowledge. Notably, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while UCE increases from 19.06 to 29.20 (10.14 ↑). A similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. UCE’s 10.29 ↑), showing the robustness of our method in erasing increasing multiple concepts.
5.2 On Multi-Concept Erasure
Evaluation setup. Another more realistic erasure scenario is multi-concept erasure, where massive concepts are required to be erased at once. Herein, we follow the experiment setup in [35] for erasing multiple celebrities, where we experiment with erasing 10, 50, and 100 celebrities and collect another 100 celebrities as non-target concepts. We prepare 5 prompt templates for each celebrity concept. For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images. For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates per concept). In evaluation, we adopt GIPHY Celebrity Detector (GCD) [20] and measure the top-1 GCD accuracy, indicated by $\text{Acc}_{e}$ for erased target concepts and $\text{Acc}_{r}$ for retained non-target concepts. Meanwhile, the harmonic mean $H_{o}$ is adopted to assess the overall erasure performance:
$H_{o}=\frac{2}{\left(1-\text{Acc}_{e}\right)^{-1}+\left(\text{Acc}_{r}\right)^{-1}}.$ (12)
Additionally, we report the results on MS-COCO to demonstrate the prior preservation of general concepts.
@@ 5.2 On Multi-Concept Erasure
Table 2: Quantitative comparison of the few-concept erasure in erasing single artistic style.

### v2

@@ 3.2 Apply Null-Space Constraints
$\left\{\mathbf{U},\mathbf{\Lambda},\mathbf{U}^{\top}\right\}=\text{SVD}\left(\mathbf{C}_{0}\mathbf{C}_{0}^{\top}\right),\ \ \mathbf{U}\in\mathbb{R}^{d_{0}\times d_{0}}\xrightarrow[\text{values}]{\text{zero singular}}\hat{\mathbf{U}}\Longrightarrow\mathbf{P}=\hat{\mathbf{U}}\hat{\mathbf{U}}^{\top}.$ (2)
The final update applied to model parameters is $\bm{\Delta}\mathbf{P}$, which projects $\bm{\Delta}$ onto the null space of $\mathbf{C}_{0}$. This ensures that updates do not interfere with non-target concepts, satisfying $\|(\bm{\Delta}\mathbf{P})\mathbf{C}_{0}\|^{2}=0$. To solve for the updates, we minimize the following objective:
$\bm{\Delta}_{\text{Null}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{\text{$e_{1}$}}\ +\ \cancel{{\color[rgb]{.5,.5,.5}\underbrace{\|\left(\bm{\Delta}\mathbf{P}\right)\mathbf{C_{0}}\|^{2}}_{e_{0}=0}}}\ +\underbrace{\|\bm{\Delta}\mathbf{P}\|^{2}}_{\text{regularization}},$ (3)
where $\|\bm{\Delta}\mathbf{P}\|^{2}$ is a regularization term to ensure convergence. The preservation term $\|(\bm{\Delta}\mathbf{P})\mathbf{C}_{0}\|^{2}$ is omitted, as it is guaranteed to be zero by the null-space constraint. This objective enables us to update the model parameters such that target concepts are effectively erased while non-target representations remain unaffected, thereby achieving prior-preserved concept erasure.
@@ 4 Prior Knowledge Refinement
However, as more diverse non-target concepts are included in the retain set, the rank of the correlation matrix $\mathbf{C}_{0}\mathbf{C}_{0}^{\top}$ increases [^2] We assume that the concepts are not exactly linearly dependent in the representation space, which is generally satisfied in practice due to the semantic diversity and high dimensionality of the embedding space.. The null space, defined as the orthogonal complement of this span, correspondingly shrinks in dimension:
$\dim(\mathrm{Null}(\mathbf{C}_{0}))=d_{0}-\mathrm{rank}(\mathbf{C}_{0}\mathbf{C}_{0}^{\top}).$ (4)
Herein, the null space dimension characterizes the degrees of freedom available for editing without affecting the retained concepts. However, as this dimension shrinks, to ensure sufficient degrees of freedom for concept erasure, we are compelled to include singular vectors w.r.t. non-zero singular values in $\hat{\mathbf{U}}$ following Fang et al. (2024), which leads to an approximate null space and induces semantic degradation within the retain set (see Fig. 2). To improve, we propose Prior Knowledge Refinement, a structured strategy for refining the retain set to enable accurate null-space construction, with three complementary techniques: Influence-Based Prior Filtering (Sec. 4.1) to discard weakly affected non-target concepts to form a viable null space; Directed Prior Augmentation (Sec. 4.2) to expand the retain set with targeted and semantically consistent variations; and Invariant Equality Constraints (Sec. 4.3) to enforce equality constraints to preserve critical invariants during generation.
@@ 4.1 Influence-Based Prior Filtering (IPF)
Given a predefined retain set, existing editing-based methods Gandikota et al. (2024); Gong et al. (2025) treat all non-target concepts equally when enforcing prior preservation. However, an overlooked fact is that parameter updates inherently induce output changes over non-target concepts, and these changes vary across different non-target concepts. This suggests that not all non-target concepts contribute equally to preserving prior knowledge, and weakly influenced concepts offer little benefit but introduce additional ranks that narrow the null space.
To this end, we propose an explicit and model-consistent metric, i.e., prior shift, to quantify how much a non-target concept is affected by concept erasure. Specifically, we isolate the effect of erasure by solving for a closed-form update $\bm{\Delta}_{\text{erase}}$ that minimizes only the erasure error $e_{1}$ while discarding the preservation term $e_{0}$ from Eq. 1:
$\bm{\Delta}_{\text{erase}}=\mathop{\text{arg\ min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\bm{\Delta}\|^{2}}_{\text{regularization}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\left(\mathbf{I}+\mathbf{C}_{1}\mathbf{C}_{1}\right)^{-1}.$ (5)
where $\|\bm{\Delta}\|^{2}$ is introduced for convergence. Then, for each non-target concept embedding $\bm{c}$, we define its prior shift as: $\|\bm{\Delta}_{\text{erase}}\bm{c}\|^{2}$. This value offers a faithful reflection of how parameter updates perturb a non-target concept in the feature space with closed-form computation, and can naturally generalize to assessing multi-concept erasure effects. Based on this, we filter the original retain set $\mathbf{R}$ to focus only on highly influenced concepts:
$\mathbf{R}_{f}:\mathbf{R}\mapsto\{\bm{c}_{0}\in\mathbf{R}\mid\|\bm{\Delta}_{\text{erase}}\bm{c}_{0}\|^{2}>\mu\},$ (6)
where the mean value $\mu=\mathbb{E}_{\bm{c}_{0}\sim\mathbf{R}}\left[\|\bm{\Delta}_{\text{erase}}\bm{c}_{0}\|^{2}\right]$ serves as a filtering threshold.
@@ 4.2 Directed Prior Augmentation (DPA)
To enhance prior preservation with broader retain coverage, an intuitive strategy is to augment the retain set by perturbing non-target embedding $\bm{c}_{0}$ with random noise Lyu et al. (2024). However, this strategy would introduce meaningless embeddings that fail to generate semantically coherent images (e.g., noise image), resulting in excessive preservation with increasing ranks. To search for more semantically consistent concepts, we introduce directed noise by projecting the random noise $\bm{\epsilon}$ onto the direction in which the model parameters $\mathbf{W}$ exhibit minimal variation. This operation ensures the perturbed embeddings express closer semantics to the original concept after being mapped by $\mathbf{W}$ in Fig. 3. Specifically, we first derive a projection matrix $\mathbf{P}_{\text{min}}$:
@@
where $\mathbf{U}_{\text{min}}={\mathbf{U}}_{\mathbf{W}}[:,-r:]$ denotes the singular vectors w.r.t. the smallest $r$ singular vectors [^3] Empirically, the model parameter matrix $\mathbf{W}$ is usually full rank, thus its all singular values are non-zero., which represent the $r$ least-changing directions of $\mathbf{W}$ and constrain the rank of the augmented embeddings to a maximum of $r$. Then the directed noise $\bm{\epsilon}\cdot\mathbf{P}_{\text{min}}$ is used to perturb the original embedding via:
$\bm{c}^{\prime}_{0}=\bm{c}_{0}+\bm{\epsilon}\cdot\mathbf{P}_{\text{min}},\quad\bm{\epsilon}\sim\mathcal{N}(\mathbf{0},\mathbf{I}).$ (8)
@@
$\mathbf{R}^{\text{aug}}:\mathbf{R}\mapsto\bigcup_{\bm{c}_{0}\in\mathbf{R}}\left\{\bm{c}_{0,k}^{\prime}\mid k=1,\dots,N_{A}\right\},$ (9)
where $N_{A}$ denotes the augmentation times and $\bm{c}_{0,k}^{\prime}$ represents the $k$-th augmented embedding given $\bm{c}_{0}\in\mathbf{R}$ using Eq. 8. In implementation, we first filter the original retain set $\mathbf{R}$ to obtain $\mathbf{R}_{f}$ using IPF. Subsequently, further augmentation and filtering are applied to $\mathbf{R}_{f}$ using DPA and IPF to obtain $\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$. Finally, we combine them to serve as the final refined retain set $\mathbf{R}_{\text{refine}}=\mathbf{R}_{f}\cup\left(\mathbf{R}_{f}\right)^{\text{aug}}_{f}$.
@@ 4.3 Invariant Equality Constraints (IEC)
In parallel, we identify certain invariants during the T2I generation process, i.e., intermediate variables that remain unchanged with varying sampling prompts. One such invariant is the CLIP-encoded [SOT] token. Since the encoding process is masked by causal attention and all prompts are prefixed with the fixed [SOT] token during tokenization, its embedding consistently remains unchanged during T2I process. Another invariant is the null-text embedding, as it corresponds to the unconditional generation under the classifier-free guidance Ho & Salimans (2022), which also remains unchanged despite prompt variations. Given the invariance of these embeddings, we consider additional protection measures to ensure their outputs remain unchanged during concept erasure. Specifically, we introduce explicit equality constraints over invariants based on Eq. 3:
$\mathop{\text{min}}\limits_{\bm{\Delta}}\underbrace{\|(\mathbf{W}+\bm{\Delta}\mathbf{P})\mathbf{C}_{1}-\mathbf{W}\mathbf{C}_{*}\|^{2}}_{e_{1}}+\underbrace{\|\bm{\Delta}\mathbf{P}\|^{2}}_{\text{regularization}},\quad\text{s.t.}\,\underbrace{(\bm{\Delta}\mathbf{P})\mathbf{C}_{2}=\mathbf{0}}_{\text{equality constraints}},$ (10)
where $\mathbf{C}_{2}$ denotes the stacked invariant embedding matrix of [SOT] and null-text. Derive the projection matrix $\mathbf{P}$ from $\mathbf{R}_{\text{refine}}$, we can compute the closed-form solution of Eq. 10 using Lagrange Multipliers from Appx. B.3:
$\left(\bm{\Delta}\mathbf{P}\right)_{\text{Ours}}=\mathbf{W}\left(\mathbf{C}_{*}\mathbf{C}_{1}^{\top}-\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\right)\mathbf{P}\mathbf{Q}\mathbf{M},$ (11)
@@
$\mathbf{M}=\left(\mathbf{C}_{1}\mathbf{C}_{1}^{\top}\mathbf{P}+\mathbf{I}\right)^{-1},\mathbf{Q}=\mathbf{I}-\mathbf{M}\mathbf{C}_{2}\left(\mathbf{C}_{2}^{\top}\mathbf{P}\mathbf{M}\mathbf{C}_{2}\right)^{-1}\mathbf{C}_{2}^{\top}\mathbf{P}.$ (12)
@@
Figure 4: Qualitative comparison of the few-concept erasure in erasing instances. The erased and preserved generations are highlighted with red and green boxes, respectively. Our method exhibits consistent prior preservation with less semantic degradation for non-target concepts. For example, the middle column better retains details such as Mickey’s hat and button count, and the right column demonstrates more consistent Hello Kitty generations along with three concepts erased.
@@ 5 Experiments
In this section, we conduct extensive experiments on three representative erasure tasks, including few-concept erasure, multi-concept erasure, and implicit concept erasure (Appx. D.4), validating our superior prior preservation. The compared baselines include ConAbl Kumari et al. (2023), MACE Lu et al. (2024a), RECE Gong et al. (2025), and UCE Gandikota et al. (2024), which have achieved SOTA performance across various concept erasure tasks. In implementation, we conduct all experiments on SDv1.4 AI (2022) and generate each image using DPM-solver sampler Lu et al. (2022) over 20 sampling steps with classifier-free guidance Ho & Salimans (2022) of 7.5. More implementation details and compared baselines can be found in Appx. C and Appx. D.3.
@@ 5.1 On Few-Concept Erasure
Evaluation setup. We evaluate few-concept erasure on instance erasure and artistic style erasure following Lyu et al. (2024), using 80 instance templates and 30 artistic style templates with 10 images per template per concept. We use two metrics for evaluation: CLIP Score (CS) Radford et al. (2021) for the text-image similarity and Fréchet Inception Distance (FID) Heusel et al. (2017) for the distributional distance before and after erasure. Following Lyu et al. (2024), we select non-target concepts with similar semantics to the target concept for comparison and report CS for targets and FID for non-targets in the main paper. Complete comparisons are presented in Appx. D.2. We further compare the generations on MS-COCO captions Lin et al. (2014), where we generate images with the first 1,000 captions, and report CS and FID to measure general knowledge preservation.
Table 1: Quantitative comparison of the few-concept erasure on instances (left) and artistic styles (right) following Lyu et al. (2024). Arrows indicate the preferred direction for each metric, and the best results are highlighted in bold. Our method consistently improves prior preservation for non-target and general concepts from MS-COCO (shaded in pink) while achieving effective concept erasure. While our CS is not the lowest for target concpet, Appx. D.1 and Fig. 7 show our method is sufficient for erasure, and lower CS may further compromise prior preservation.
@@
Analysis and discussion. Table 1 compares the results of erasing various instance concepts and artistic styles. Our method consistently achieves the lowest FIDs across all non-target concepts, demonstrating superior prior preservation with minimal alteration to the original content. Moreover, we emphasize that our erasure is sufficiently effective, even without achieving the lowest CS, as shown in Fig. 4 and Appx. D.1. In contrast, lower CS values typically indicate over-erasure, which results in excessive degradation of prior knowledge. Notably, with the number of target concepts increasing from 1 to 3, our FID in Pikachu rises from 16.81 to 21.40 (4.59 ↑), while UCE increases from 19.06 to 29.20 (10.14 ↑). A similar pattern is observed in Hello Kitty (Our 4.48 ↑ v.s. UCE’s 10.29 ↑), showing our robustness in erasing increasing target concepts.
5.2 On Multi-Concept Erasure
Evaluation setup. Another more realistic erasure scenario is multi-concept erasure, where massive concepts are required to be erased at once. We follow the experiment setup in Lu et al. (2024a) for erasing multiple celebrities, where we experiment with erasing 10, 50, and 100 celebrities and collect another 100 celebrities as non-target concepts. Specifically, we prepare 5 prompt templates for each celebrity concept. For non-target concepts, we generate 1 image per template for each of the 100 concepts, totaling 500 images. For target concepts, we adjust the per-concept quantity to maintain a total of 500 images (e.g., erasing 10 celebrities involves generating 10 images with 5 templates per concept). In evaluation, we adopt GIPHY Celebrity Detector (GCD) Hasty et al. (2019) and measure the top-1 GCD accuracy, indicated by $\text{Acc}_{e}$ for erased target concepts and $\text{Acc}_{r}$ for retained non-target concepts. Meanwhile, the harmonic mean $H_{o}=\frac{2}{\left(1-\text{Acc}_{e}\right)^{-1}+\left(\text{Acc}_{r}\right)^{-1}}$ is adopted to assess the overall erasure performance. Additionally, we report the results on MS-COCO to demonstrate the prior preservation of general concepts.
Figure 5: Quantitative comparison of the multi-concept erasure in erasing celebrities (celeb). The erased and preserved generations are marked with red and green boxes. Our method precisely erases 100 celebrities while preserving generations of other non-target concepts.
Table 2: Quantitative comparison of the multi-concept erasure in erasing 10, 50, and 100 celebrities. The best results are highlighted in bold. Our method can effectively erase up to 100 celebrities simultaneously, achieving low $\text{Acc}_{e}$ (%) and high $\text{Acc}_{r}$ (%) that preserve non-target celebrities with minimal appearance changes. This yields the best overall erasure performance $H_{o}$ and competitive runtime (s) on one A100 GPU, successfully erasing 100 concepts in just 5 seconds.